# Down with the Hierarchy: The ‘H’ in HNSW Stands for “Hubs”

### v1

@@ Abstract.
Driven by recent breakthrough advances in neural representation learning, approximate near-neighbor (ANN) search over vector embeddings has emerged as a critical computational workload. With the introduction of the seminal Hierarchical Navigable Small World (HNSW) algorithm, graph-based indexes have established themseves as the overwhelmingly dominant paradigm for efficient and scalable ANN search. As the name suggests, HNSW searches a layered hierarchical graph to quickly identify neighborhoods of similar points to a given query vector. But is this hierarchy even necessary? A rigorous experimental analysis to answer this question would provide valuable insights into the nature of algorithm design for ANN search and motivate directions for future work in this increasingly crucial domain. To that end, we conduct an extensive benchmarking study covering more large-scale datasets than prior investigations of this question. We ultimately find that a flat graph retains all of the benefits of HNSW on high-dimensional datasets, with latency and recall performance essentially identical to the original algorithm but with less memory overhead. Furthermore, we go a step further and study why the hierarchy of HNSW provides no benefit in high dimensions, hypothesizing that navigable small world graphs contain a well-connected, frequently traversed “highway” of hub nodes that maintain the same purported function as the hierarchical layers. We present compelling empirical evidence that the Hub Highway Hypothesis holds for real datasets and investigate the mechanisms by which the highway forms. The implications of this hypothesis may also provide future research directions in developing enhancements to graph-based ANN search.
@@ 1. Introduction
Near neighbor search is a fundamental problem in computational geometry that lies at the heart of countless practical applications. From industrial-scale recommendation (Feng et al., 2022) to retrieval-augmented generation (Lewis et al., 2020) and even to computational biology (Zhao et al., 2024), numerous data-intensive tasks utilize similarity search at some location in the stack. As a result, similarity indexes are very well-studied (Guo et al., 2019; Malkov and Yashunin, 2016; Johnson et al., 2017; Aguerrebere et al., 2023; Subramanya et al., 2019) with multiple large-scale benchmarks and leaderboards to compare techniques (Aumüller et al., 2018; Simhadri et al., 2022).
@@
Historically, the state-of-the art for near neighbor search involved constructing sophisticated tree-based data structures, such as $kd$-trees (Bentley, 1975) and cover trees (Beygelzimer et al., 2006), that guaranteed exact solutions while avoiding a brute-force examination of all points. However, the recent advent of large-scale neural representation learning, including large language models (LLMs), placed a significant strain on these classical methods that were developed to target a much lower-dimensional search space. In response, the community has turned to approximate search methods. While alternative approximate indexing methods such as locality-sensitive hashing (Indyk and Motwani, 1998) and product quantization (Jegou et al., 2010), have garnered significant interest, graph-based approaches generally achieve the strongest performance on established ANN benchmarks (Aumüller et al., 2018; Simhadri et al., 2022). Introduced in 2016, the Hierarchical Navigable Small World (HNSW) algorithm (Malkov and Yashunin, 2016), emerged as one of the first high-performance graph-based search indexes at scale and still enjoys immense popularity to this day with over 4300 Github stars [^1] https://github.com/nmslib/hnswlib.
@@
As the name implies, a core feature of the HNSW index is its hierarchically layered graph akin to a skip list (Pugh, 1990) where a search iteratively traverses through graphs of increasing density before converging to a neighborhood of similar points in the final graph layer. By drawing intuition from skip lists, the HNSW authors argue that the initial coarse graph layers allow for efficiently identifying the neighborhood of similar points in the collection through fewer overall comparisons.
@@
Despite the immense popularity of HNSW, however, the algorithm suffers from multiple scalability bottlenecks. For instance, the hierarchical layers introduce a hefty memory overhead. Moreover, Malkov and Yashunin (2016) also note that the hierarchy can reduce the overall system throughput in distributed settings when compared to a flat graph.
@@
While the high cost of this hierarchy in HNSW is traditionally justified by the low latency of graph-based indexes, we ask the question of whether the hierarchy is even necessary in the first place. A growing body of results suggest that the hierarchy may be a vestigial artifact left behind by the lower-dimensional problems of the past. For example, Lin and Zhao (2019) published a set of experiments suggesting that the hierarchical component of HNSW is only beneficial for low dimensional inputs ($d<32$). Other recent works have observed similar problems with the hierarchy. For example, Coleman et al. (2022) present an ablation study showing the same behavior in the appendix of their paper. This behavior seems to be an increasingly well-known “folklore” of the similarity search community, but there is currently an absence of a thorough investigation and analysis of this phenomenon.
@@
An exhaustive benchmark comparison on the utility of the HNSW hierarchy would go a long way towards improving our understanding of graph-based similarity search in high-dimensional space and inform future research directions in developing improvements. However, there are several key difficulties with executing such a benchmarking study. The first issue is that running a meaningful ablation study requires first engineering a similarity search implementation over a flat graph that reaches performance parity with HNSW and other state of the art approaches. This is an increasingly nontrivial task, as the community has invested significant performance engineering efforts into these codebases such that a non-expert implementation has little hope of being competitive.
@@
The second challenge is designing benchmarking experiments in a manner that avoids the confounding effect of the performance of particular software implementation when making conclusions about an algorithm’s efficacy. Finally, benchmarking near-neighbor search can be highly nuanced. For example, it is possible that hierarchical structures are strongly beneficial for isolated slices of the input. These may sharply reduce the 99th percentile latency while not affecting the median.
@@
Figure 1. We hypothesize that in high dimensions, graph-based $k$-NN indices naturally form a “highway-feeder” structure, where a small subset of nodes and edges are easily reached, well-connected, and heavily traversed.
@@ 1.1. Contributions
In this paper, we take on the question of whether the hierarchical component of HNSW is truly necessary, or if we can achieve the same performance on large-scale benchmarks with simply a flat navigable small world graph. In particular, we make the following contributions:
@@
Why does hierarchy not help? We hypothesize that the hierarchy benefits decrease due to hubness. Hubness is a high-dimensional phenomenon that causes a skewed distribution in the near-neighbor lists of search queries. We hypothesize that hubness leads to preferential attachment in the similarity search graph, inducing the formation of easily-traversed highways that connect disparate regions of the graph. This hypothesis, which we call the hub-highway hypothesis, explains why we no longer need the hierarchy in high dimensions; we can simply traverse the intrinsic highway structure that naturally forms in high-dimensional spaces. To that end, we conduct a series of experiments to investigate the hub-highway hypothesis. Our results ultimately support this hypothesis, showing that hubness is responsible for driving the connectivity of similarity search graphs. This insight opens up exciting new research directions in graph construction, link pruning, and graph traversal. Our work also reveals that the hub-highway is an intrinsic and naturally-forming structure in high-dimensional proximity graphs, which we believe has important implications for the scaling potential of many traversal heuristics. [^2] Specifically, that the benefits of sophisticated initialization methods will decay under hub-style preferential attachment.
@@
Takeaway: Our benchmarks reveal that HNSW can be significantly optimized for modern high-dimensional embedding workloads. In particular, our implementation saves roughly $38\%$ of peak memory consumption compared to hnswlib (and sizable further headroom is likely). Our results confirm the folklore of the similarity search community, conclusively demonstrating that we can remove the hierarchy on high-dimensional inputs with impunity.
@@ 2. Background: Similarity Search & HNSW
where $\phi$ is usually the $\ell_{2}$ or cosine distance. With the enormity of modern data workloads and the underlying vector dimensionality, it becomes computationally infeasible to exhaustively search for the true top-$k$ neighbors for any query $q.$ Thus, approximate search algorithms trade-off quality of the search for lower latency.
@@ 2.1. HNSW Overview
With this formalization of the ANNS problem, we will now briefly review the key elements of the HNSW algorithm, which is the central focus of our benchmarking study. As we alluded to previously, HNSW builds off of prior work in navigable small world graph indexes which was introduced in (Malkov et al., 2014). Small world graphs are a well-studied phenomenon in both computing and the social sciences and are primarily defined by the fact that the average length of a shortest path between two vertices is small (typically scaling logarithmically with the number of nodes in the network) (Travers and Milgram, 1977; Watts and Strogatz, 1998; Kleinberg, 2000). Small world graphs are also often characterized by the presence of well-connected hub nodes which we discuss further in the next section.
@@
Specifically, the HNSW index is constructed in an iterative fashion. For a newly inserted element $x$, the algorithm will randomly select a maximum layer $l$ and then insert the new point into every layer up to $l$. This randomized process is executed with an exponentially decaying probability distribution such that, in expectation, each subsequent layer has exponentially more nodes than its predecessor. Within a layer, HNSW greedily adds edges between $x$ and its $ef$ closest neighbors (where $ef$ is a hyperparameter) where the neighbors consist of previously inserted points. This process then repeats in the subsequent layer below using the closest neighbors found in the prior graph as entry points. Through this process, the top layer of the hierarchy will be the coarsest directed graph, consisting of the fewest nodes and edges, and the bottom layer will be the densest and contain all of the nodes, each with connections to $ef$ number of neighbors. As an additional, and important, optimization, HNSW also implements the pruning heuristic of (Arya and Mount, 1993) that will prune an edge from $u$ to $v$ if there exists another edge from $u$ to a neighbor $w$ of $v$ such that the distance from $u$ to $w$ is less than that of $u$ to $v$.
The search procedure of HNSW also executes iteratively where the algorithm maintains a list of candidate points at each layer of the hierarchical graph before returning the final list $k$ nearest neighbor candidates after traversing the base graph layer.
@@ 3.1. Near-Neighbor Benchmarks
Due to the importance of the $k$-NNS problem, there have been several large-scale near-neighbor benchmarks in recent years. The original ANN Benchmarks  (Aumüller et al., 2018) made significant impact because it provided the first standard evaluation of near-neighbor search algorithms. Since its inception, the benchmark has grown in scope from a handful of algorithms to include over 30 methods on 9 datasets. However, most of the datasets included in ANN Benchmarks are relatively small by modern standards at around one million points. To reflect the growing need for high-dimensional, large-scale embedding search, NeurIPS hosted a competition track known as Big ANN Benchmarks (Simhadri et al., 2022), which evaluates algorithms on five billion-scale datasets. Graph-based indexes easily and consistently place in among the top three algorithms in each benchmark. However, these benchmarks are not sufficient for our purposes because they focus exclusively on overall throughput metrics and do not report tail latency numbers such as the 99th percentile, which is important in our case for teasing apart the impact of hierarchy in HNSW.
Hierarchy Studies: Hierarchical structures are considered a fundamental component of graph-based near-neighbor algorithms. As a result, numerous high-performing algorithms use some form of hierarchy including HNSW (Malkov and Yashunin, 2016), ONNG (Iwasaki and Miyazaki, 2018) PANNG (Iwasaki, 2016), and HCNNG (Munoz et al., 2019). However, this practice has recently come under scrutiny. Dobson et al. (2023) show that the hierarchy may be unnecessary for some workloads, based on the observation that HNSW under-performs both HCNNG (which uses a shallower hierarchy) and DiskANN (Subramanya et al., 2019) (which does not use a hierarchy). Lin and Zhao (2019) present results on synthetic data showing that the hierarchy is only beneficial for low-dimensional data ($d<32$), but analyze only a limited number of small real-world datasets and – most importantly – do not explain why the hierarchy fails to provide value. In this work, we aim to address these gaps in the literature through more comprehensive benchmarks coupled with an analysis of why hierarchical structures in high-dimensional space may not add value. In the following section, we aim to reproduce the results of both Malkov and Yashunin (2016) and Lin and Zhao (2019) to confirm that we can independently replicate their findings with our own software implementation before proceeding to new experiments on larger-scale benchmark datasets.

### v2

@@ Abstract.
Driven by recent breakthrough advances in neural representation learning, approximate near-neighbor (ANN) search over vector embeddings has emerged as a critical computational workload. With the introduction of the seminal Hierarchical Navigable Small World (HNSW) algorithm, graph-based indexes have established themselves as the overwhelmingly dominant paradigm for efficient and scalable ANN search. As the name suggests, HNSW searches a layered hierarchical graph to quickly identify neighborhoods of similar points to a given query vector. But is this hierarchy even necessary? A rigorous experimental analysis to answer this question would provide valuable insights into the nature of algorithm design for ANN search and motivate directions for future work in this increasingly crucial domain. To that end, we conduct an extensive benchmarking study covering more large-scale datasets than prior investigations of this question. We ultimately find that a flat navigable small world graph graph retains all of the benefits of HNSW on high-dimensional datasets, with latency and recall performance essentially identical to the original algorithm but with less memory overhead. Furthermore, we go a step further and study why the hierarchy of HNSW provides no benefit in high dimensions, hypothesizing that navigable small world graphs contain a well-connected, frequently traversed “highway” of hub nodes that maintain the same purported function as the hierarchical layers. We present compelling empirical evidence that the Hub Highway Hypothesis holds for real datasets and investigate the mechanisms by which the highway forms. The implications of this hypothesis may also provide future research directions in developing enhancements to graph-based ANN search.
@@ 1. Introduction
Near neighbor search is a fundamental problem in computational geometry that lies at the heart of countless practical applications. From industrial-scale recommendation (Feng et al., 2022) to retrieval-augmented generation (Lewis et al., 2020) and even to computational biology (Zhao et al., 2024), numerous data-intensive tasks utilize similarity search at some location in the stack. As a result, similarity indexes are very well-studied (Guo et al., 2019; Malkov and Yashunin, 2016; Johnson et al., 2017; Aguerrebere et al., 2023; Jayaram Subramanya et al., 2019) with multiple large-scale benchmarks and leaderboards to compare techniques (Aumüller et al., 2018; Simhadri et al., 2022).
@@
Historically, the state-of-the-art for near neighbor search involved constructing sophisticated tree-based data structures, such as $kd$-trees (Bentley, 1975) and cover trees (Beygelzimer et al., 2006), that guaranteed exact solutions while avoiding a brute-force examination of all points. However, the recent advent of large-scale neural representation learning, including large language models (LLMs), places a significant strain on these classical methods that were developed to target a much lower-dimensional search space. In response, the community has turned to approximate search methods. While alternative approximate indexing methods such as locality-sensitive hashing (Indyk and Motwani, 1998) and product quantization (Jegou et al., 2010), have garnered significant interest, graph-based approaches generally achieve the strongest performance on established ANN benchmarks (Aumüller et al., 2018; Simhadri et al., 2022). Introduced in 2016, the Hierarchical Navigable Small World (HNSW) algorithm (Malkov and Yashunin, 2016), emerged as one of the first high-performance graph-based search indexes at scale and still enjoys immense popularity to this day with over 4300 Github stars [^1] https://github.com/nmslib/hnswlib.
@@
As the name implies, a core feature of the HNSW index is its hierarchically layered graph akin to a skip list (Pugh, 1990) where the search process iteratively traverses through graphs of increasing density before converging to a neighborhood of similar points in the final graph layer. By drawing intuition from skip lists, the HNSW authors argue that the initial coarse graph layers allow for efficiently identifying the neighborhood of similar points in the collection through fewer overall comparisons.
@@
Despite the immense popularity of HNSW, however, the algorithm suffers from multiple scalability bottlenecks. For instance, the hierarchical layers introduce a hefty memory overhead. Moreover, Malkov and Yashunin (2016) also note in Section 6 of their paper that the hierarchy can reduce the overall system throughput in distributed settings when compared to a flat NSW graph.
@@
While the high cost of this hierarchy in HNSW is traditionally justified by the low latency of graph-based indexes, we ask the question of whether the hierarchy is even necessary in the first place. A growing body of results suggest that the hierarchy may be a vestigial artifact left behind by the lower-dimensional problems of the past. For example, Lin and Zhao (2019) published a set of experiments suggesting that the hierarchical component of HNSW is only beneficial for low-dimensional inputs ($d<32$). Other recent works have observed similar problems with the hierarchy. For example, Coleman et al. (2022) present an ablation study showing the same behavior in the appendix of their paper. This behavior seems to be an increasingly well-known “folklore” of the similarity search community, but there is currently an absence of a thorough investigation and analysis of this phenomenon.
@@
An exhaustive benchmark comparison on the utility of the HNSW hierarchy would go a long way towards improving our understanding of graph-based similarity search in high-dimensional space and inform future research directions in developing improvements. However, there are several key difficulties with executing such a benchmarking study. The first issue is that running a meaningful ablation study requires first engineering a similarity search implementation over a flat NSW graph that reaches performance parity with HNSW and other state of the art approaches (Douze et al., 2024; Sun et al., 2024; Jayaram Subramanya et al., 2019). This is an increasingly nontrivial task, as the community has invested significant performance engineering efforts into these codebases such that a non-expert implementation has little hope of being competitive.
@@
The second challenge is designing benchmarking experiments in a manner that avoids the confounding effect of the performance of a particular software implementation when making conclusions about the efficacy of the algorithm. Finally, benchmarking near-neighbor search can be highly nuanced. For example, it is possible that hierarchical structures are strongly beneficial for isolated slices of the input. These may sharply reduce the 99th percentile latency while not affecting the median.
@@
Figure 1. We hypothesize that in high dimensions, graph-based ANN indexes naturally form a “highway-feeder” structure, where a small subset of nodes and edges are easily reached, well-connected, and heavily traversed.
@@ 1.1. Contributions
In this paper, we address the question of whether the hierarchical component of HNSW is truly necessary. Our central research question is, “Can we achieve the same performance on large-scale benchmarks with simply a flat navigable small world graph?” In summary, we make the following contributions:
@@
Why does hierarchy not help? We hypothesize that the hierarchy benefits decrease in high-dimensions due to hubness. Hubness is a high-dimensional phenomenon that causes a skewed distribution in the near-neighbor lists of search queries (Radovanovic et al., 2010). We hypothesize that hubness leads to preferential attachment in the similarity search graph, inducing the formation of easily-traversed highways that connect disparate regions of the graph. This hypothesis, which we call the Hub Highway Hypothesis, explains why we no longer need the hierarchy in high dimensions; we can simply traverse the intrinsic highway structure that naturally forms in high-dimensional spaces. To that end, we conduct a series of experiments to investigate the hub-highway hypothesis. Our results ultimately support this hypothesis, showing that hubness is responsible for driving the connectivity of similarity search graphs. This insight opens up exciting new research directions in graph construction, link pruning, and graph traversal.
@@
Practical implications: Our benchmarks reveal that HNSW can be significantly optimized for modern high-dimensional embedding workloads. For instance, our implementation saves roughly $38\%$ and $39\%$ of peak memory consumption on two Big-ANN benchmark datasets compared to hnswlib (and sizable further headroom is likely). Our results confirm the folklore of the similarity search community, conclusively demonstrating that we can remove the hierarchy on high-dimensional inputs with impunity.
@@ 2. Background: Similarity Search & HNSW
where $\phi$ is usually the $\ell_{2}$ or cosine similarity. With the enormity of modern data workloads and the underlying vector dimensionality, it becomes computationally infeasible to exhaustively search for the true top-$k$ neighbors for any query $q.$ Thus, approximate search algorithms trade-off quality of the search for lower latency.
@@ 2.1. HNSW Overview
With this formalization of the ANNS problem, we will now briefly review the key elements of the HNSW algorithm, which is the central focus of our benchmarking study. As we alluded to previously, HNSW builds off of prior work in navigable small world graph indexes introduced in (Malkov et al., 2014). Small world graphs are a well-studied phenomenon in both computing and the social sciences and are primarily defined by the fact that the average length of a shortest path between two vertices is small (typically scaling logarithmically with the number of nodes in the network) (Travers and Milgram, 1977; Watts and Strogatz, 1998; Kleinberg, 2000). Small world graphs are also often characterized by the presence of well-connected hub nodes which we discuss further in the next section.
Algorithm 1 HNSW Construction
1:Input: Set of data points $D$, max layer $L_{max}$, max connections per layer $M$, layer insertion probability $m_{l}$, size of dynamic candidate list $efc$
2:Output: HNSW graph with hierarchical layers
3:procedure Construct($D,L_{max},M,m_{l},efc$)
4:     Initialize empty hierarchical graph $G$
5:     Initialize entry point $ep\leftarrow\text{None}$
6:     for each $p\in D$ do
7:         $L_{p}\leftarrow$ GeometricDistribution($m_{l}$)
8:         if $ep=\text{None}$ then
9:              Set $p$ as entry point $ep$
10:              Insert $p$ into all levels $\leq L_{p}$
11:         end if
12:         for $l=L_{max}$ to $L_{p}$ do
13:              $ep\leftarrow$ SearchLayer($G,l,p,ep,efc$) $\triangleright$ Algorithm 2
14:         end for
15:         for $l=0$ to $L_{p}$ do
16:              $N\leftarrow$ SelectNeighbors($p,G,l,M$)
17:              Add edges from $q$ to each neighbor $n\in N$ at layer $l$
18:              if $n\in N$ has $<M$ edges then
19:                  Add back-connections to $q$ to node $n$.
20:              else
21:                  Run SelectNeighbors on $\{q,$ edges of $n\}$.
22:              end if
23:         end for
24:     end for
25:     if $L_{p}>L_{ep}$ then
26:         $ep\leftarrow p$
27:     end if
28:end procedure
29:
30:function SelectNeighbors($p,G,l,M$)
31:     Compute distances from $p$ to all nodes in $G[l]$
32:     Return $M$ nodes based on selection heuristic in  (Arya and Mount, 1993)
33:end function
2.2. HNSW Search Algorithm
Algorithm 2 HNSW Query
1:Input: Graph $G$, layer $l$, query $q$, starting point $p$, number of nearest neighbors to return $efs$
2:procedure SearchLayer($G,l,q,p,efs$)
3:     Candidate queue $C={p}$, currently top results queue $T={p}$, visited list $V={p}$
4:     while $C$ is not empty do
5:         $c\leftarrow\text{nearest element from}~{}C~{}\text{to}~{}q$
6:         $f\leftarrow\text{furthest element from}~{}T~{}\text{to}~{}q$
7:         if dist(c, q) ¿ dist(f, q) then return T
8:         end if
9:         for $e\in\text{neighbourhood}(c)~{}\text{at layer}~{}l$ do
10:              if $e\in V$ then
11:                  continue
12:              end if
13:              $V.add(e)$
14:              if $dist(e,q)\leq dist(f,q)~{}\text{or}~{}|T|\leq efs$ then
15:                  $C.add(e)$
16:                  $T.add(e)$
17:              end if
18:              if $|T|\geq efs$ then
19:                  Remove furthest point to q from T
20:              end if
21:              $f\leftarrow\text{furthest element from}~{}T~{}\text{to}~{}q$
22:         end forreturn T
23:     end while
24:end procedure
@@ 2.2. HNSW Search Algorithm
Specifically, the HNSW index is constructed in an iterative fashion. For a newly inserted element $x$, the algorithm will randomly select a maximum layer $l$ and then insert the new point into every layer up to $l$. This randomized process is executed with an exponentially decaying probability distribution such that, in expectation, each subsequent layer has exponentially more nodes than its predecessor. Within a layer, HNSW greedily adds edges between $x$ and its $M$ closest neighbors (where $M$ is a hyperparameter) where the neighbors consist of previously inserted points. This process then repeats in the subsequent layer below using the closest neighbors found in the prior graph as entry points. Through this process, the top layer of the hierarchy will be the coarsest directed graph, consisting of the fewest nodes and edges, and the bottom layer will be the densest and contain all of the nodes, each with connections to (up to) $M$ neighbors. As an additional, and important, optimization, HNSW also implements the pruning heuristic of (Arya and Mount, 1993) that will prune an edge from $u$ to $v$ if there exists another edge from $u$ to a neighbor $w$ of $v$ such that the distance from $u$ to $w$ is less than that of $u$ to $v$.
The search procedure of HNSW, described in Algorithm 2 also executes iteratively where the algorithm maintains a list of candidate points at each layer of the hierarchical graph before returning the final list $k$ nearest neighbor candidates after traversing the base graph layer.
Figure 2. Visualization of the HNSW search index. We benchmark the full hierarchical graph search procedure against simply searching in the bottom base layer.
@@ 3.1. Near-Neighbor Benchmarks
Due to the importance of the $k$-NNS problem, there have been several large-scale near-neighbor benchmarks in recent years. The original ANN Benchmarks  (Aumüller et al., 2018) made significant impact because it provided the first standard evaluation of near-neigh3bor search algorithms. Since its inception, the benchmark has grown in scope from a handful of algorithms to include over 30 methods on 9 datasets. However, most of the datasets included in ANN Benchmarks are relatively small by modern standards at around one million points. To reflect the growing need for high-dimensional, large-scale embedding search, NeurIPS hosted a competition track known as Big ANN Benchmarks (Simhadri et al., 2022), which evaluates algorithms on five billion-scale datasets. Graph-based indexes easily and consistently place among the top three algorithms in each benchmark. However, these benchmarks are not sufficient for our purposes because they focus exclusively on overall throughput metrics and do not report tail latency numbers such as the 99th percentile, which is important in our case for teasing apart the impact of hierarchy in HNSW.
Hierarchy Studies: Hierarchical structures are considered a fundamental component of graph-based near-neighbor algorithms. As a result, numerous high-performing algorithms use some form of hierarchy including HNSW (Malkov and Yashunin, 2016), ONNG (Iwasaki and Miyazaki, 2018) PANNG (Iwasaki, 2016), and HCNNG (Munoz et al., 2019). However, this practice has recently come under scrutiny. Dobson et al. (2023) show that the hierarchy may be unnecessary for some workloads, based on the observation that HNSW under-performs both HCNNG (which uses a shallower hierarchy) and DiskANN (Jayaram Subramanya et al., 2019) (which does not use a hierarchy). Lin and Zhao (2019) present results on synthetic data showing that the hierarchy is only beneficial for low-dimensional data ($d<32$), but analyze only a limited number of small real-world datasets and – most importantly – do not explain why the hierarchy fails to provide value. In this work, we aim to address these gaps in the literature through more comprehensive benchmarks coupled with an analysis of why hierarchical structures in high-dimensional space may not add value. In the following section, we aim to reproduce the results of both Malkov and Yashunin (2016) and Lin and Zhao (2019) to confirm that we can independently replicate their findings with our own software implementation before proceeding to new experiments on larger-scale benchmark datasets.